# footnotetext: Corresponding author

### v3

@@ 1 Introduction
State space models (SSMs) have recently shown significant potential in addressing challenges related to modeling sequences with low latency [12]. Inspired by continuous state space models from control systems and enhanced by HiPPO initialization [13], SSMs [16] show promise in addressing long-term forgetting issue. These advancements have been integrated into large-scale representation models [38, 41]. Some pioneering works have applied SSMs for tasks like language understanding [38, 41], content-based reasoning [12], and visual recognition [35, 57]. In our work, we further explore the potential of SSMs in co-speech gesture synthesis. We observe that directly applying the selective scan mechanism from Mamba [12] to gesture generation as a sequence modeling model would result in jittery outputs. To refine the generated gestures, we propose a two-stage modeling strategy. In the first training stage, we enhance the discrete motion priors derived from VQVAEs [53] by integrating velocity and acceleration losses. In the second stage, by utilizing motion priors from VQVAEs, we introduce individual learnable queries for different body parts, thereby alleviating the jittering issue. Meanwhile, considering that the direct application of Mamba encounters the challenge of limb movements across different body parts tending to average out, we propose a hybrid scanning approach in the second stage to enhance the motion representation in the latent space. Specifically, we refine the design of spatial and temporal modeling within latent spaces by introducing a local-to-global modeling strategy and integrating attention mechanisms along with a selective scanning approach into the framework’s design. Considering the significant differences in deformation and movement patterns among different body parts [32], we propose local and global scan modules for refining the latent space representations of the movements across various body parts. These approaches enable dynamic interaction and iterative refinement of different body parts while maintaining low latency, leading to more diverse and rhythmic gestures. Our contributions can be summarized as below:
@@ 3.1 Preliminaries
where $(\Delta A)^{-1}$ denotes the inverse of matrix $\Delta A$, $I$ denotes the identity matrix. The scan module within Mamba is designed to capture temporal patterns and dependencies across multiple time steps by applying a set of trainable parameters or operations to each segment of the input sequence. In our framework, Mamba serves as a sequence modeling tool for decoding gesture actions across different parts of the body. By modifying the decoder’s input and the range of features, we leverage Mamba to model local body features and global characteristics separately. These operations are learned during training and help the model to process sequential data.
@@ 3.3 Speech-Driven Selective State Spaces Gesture Synthesis Model
Speech Feature Extraction. For audio feature extraction, two CNN-based audio feature extraction networks are employed to respectively extract amplitude and onset features. Specifically, we integrate amplitude and onset features along the channel dimension to obtain audio features $f_{A}=\{fa_{1},...,fa_{N}\}$. For processing speech input words, we employ pre-trained FastText [5] to obtain word embeddings, which are then refined by linear projections to produce text features $f_{T}=\{ft_{1},...,ft_{N}\}$. We further fuses features from the input modalities (e.g., audio and text features). The speaker embeddings $s_{id}$ are first combined with audio and text features through additive operation. By concatenating feature vectors along the channel dimension, we further apply linear transformations to determine the weight factors, and then integrating the features through an element-wise summation. The process can be formalized as:
@@
Local and Global Scans. Recognizing the diverse deformations and motion patterns in various body parts, we propose using multiple local scan modules to model the movements of different body parts (e.g., face, hand, upper and lower body) with fused multi-modal features from previous modules. To enhance the generalization of the model, we have incorporated additional learnable queries to foster the queries’ ability to learn motion patterns. As shown in Figure 2, the queries ($Q$) are integrated with the input’s multi-modal features through a multi-head cross attention mechanism ($\mathcal{F}_{\text{MHCA}}$). This allows the queries to learn the most relevant information from the speech input. The process can be formally defined as follows:
$f_{\text{face}}=\mathcal{F}_{\text{MHCA}}(Q_{\text{face}},\bar{f}_{T},\bar{f}_{A}),$ (7)
where $f_{\text{face}}$ denotes the feature of facial motion. This formulation can be applied similarly to obtain $f_{hand}$, $f_{upper}$, and $f_{lower}$ by substituting the corresponding query vectors $Q_{hand}$, $Q_{upper}$, and $Q_{lower}$ respectively. Utilizing the extracted perceptual features from various body parts, we proceed to employ Mamba to extract temporal features from the sequence, which can be formalized as belows:
${F}_{face}$ $=\text{Mamba}({f}_{face}),$ (8)
where $F_{\text{face}}$ corresponds to the temporal features of facial motion. The same approach is applied to generate temporal features for the hand, upper body, and lower body by inputting their respective perceptual features $f_{hand}$, $f_{upper}$, and $f_{lower}$ into the corresponding Mamba modules.
@@
By acquiring the distinctive features of various body parts, we improve the perception of motion patterns among them using a global scan module. Initially, we combine the features along the sequence dimension. Subsequently, by utilizing self-attention mechanism($\mathcal{F}_{\text{MHSA}}$), we model the global information across different sequence tokens. These refined features are fed into Mamba to extract temporal perceptual information, which can be fomulized as below:
@@
$\begin{gathered}F_{parts}=[{F}_{face},{F}_{hand},{F}_{upper},{F}_{lower}]\\ {f}_{mixed}=\mathcal{F}_{\text{MHSA}}(F_{parts}),\\ {f}_{global}=\text{Mamba}({f}_{mixed}),\\ \end{gathered}$ (9)
where [][][ ] denotes the concatenation operation of ${F}_{face}$, ${F}_{hand}$, ${F}_{upper}$ and ${F}_{lower}$ in the dimension of the sequence. The extracted latent features are then fed into their respective VQ-Decoders to produce the final motion predictions.
@@
Training Objectives. The model’s training objectives are composed by a composite loss function that harmonizes reconstruction and cross-entropy losses. This design aims to augment the accuracy of motion generation, encompassing the face, hands, upper and lower body. The latent reconstruction loss, represented by $L_{reclatent}$, is quantified using the Mean Squared Loss (MSELoss). Here, $z_{i}$ corresponds to the true latent vectors, while $\hat{z}_{i}$ are the vectors reconstructed by the model. The latent reconstruction loss is expressed as:
@@
where $N$ denotes the number of frames. Concurrently, to encourage diversity in the generated motions, we optimize the cross-entropy loss for latent code class classification $L_{cls}$. Specifically, we employ the Negative Log Likelihood Loss (NLLLoss), where $y_{i}$ represents the true class labels for each sample, and $\hat{y}_{i}$ denotes the model’s predicted class labels. This loss is calculated as the negative sum of the logarithm of the predicted probabilities for the correct classes:
@@ 4.1 Experiments Setting
We train and evaluate on the BEATX-standard dataset proposed by [32]. BEATX-standard contains 60 hours of data with high finger quality for 25 speakers (12 female and 13 male). The dataset comprises 1762 sequences, each with an average duration of 65.66 seconds. Each sequence includes a response to a daily inquiry. We split datasets into 85%/7.5%/7.5% for the train/val/test set.
@@ 4.4 Quantitative Results
As shown in Table 1, our method attains the highest BC, with a significant increase of 14.3% when compared to the previously top-performing method, which underscores the superior capability of MambaTalk in discerning and correlating audio-motion beats. Furthermore, our method performs competitively with EMAGE in terms of FGD, and much better than others. This emphasizes the high quality and naturalness of our generated movements, showing the ability of MambaTalk to capture real motion dynamics. This also demonstrates the authenticity of our generated motions, affirming the successful capture of inherent motion characteristics. Some results are marked as “-” because these methods can not generate facial movements. Moreover, our method outperforms previous methods in terms of MSE and LVD, with substantial improvements of 36% and 14%, respectively. These two enhancements highlight the superior accuracy and fidelity of our method in capturing fine-grained details, affirming its efficacy in synthesizing realistic and authentic facial motions.
Table 1: Quantitative results on BEATX. FGD (Frechet Gesture Distance) multiplied by $10^{-1}$, BC (Beat Constancy) multiplied by $10^{-1}$, Diversity, MSE (Mean Squared Error) multiplied by $10^{-7}$, and LVD (Learned Vector Distance) multiplied by $10^{-5}$. The best results are in bold.
@@
MambaTalk (Ours) NeurlPS 2024 5.409 $\mathbf{7.930}$ 13.07 $\mathbf{7.512}$ $\mathbf{7.496}$
@@ 4.5 Qualitative Analysis
User Study. We conducted a user study to assess the visual quality of the generated co-speech 3D gestures. For each method under comparison, we produced 10 gesture samples, which were then converted into video clips for evaluation by 39 participants. In each evaluation session, participants were presented with 20 seconds video clips generated by various models. They were instructed to assess the clips across the following dimensions: (i) naturalness, (ii) appropriateness, (iii) synchrony and (iv) smoothness. For naturalness, they evaluated the similarity of the generated gestures to those made by humans, paying attention to the authenticity and smoothness of the movements. In terms of appropriateness, they considered the alignment of the gestures with the spoken content, taking into account both the explicit meaning and the underlying semantics. For synchrony assessment, they examined the timing of the gestures in relation to the speech rhythm, audio, and facial expressions to ensure a harmonious and integrated performance. For smoothness, they assessed the gestures for any abrupt stops or unnatural jerks that might indicate a lack of fluidity in motion. We mainly compared two state-of-art methods with our proposed method (with and without VQVAE): CaMN [33], EMAGE [32], and the ground truth. As presented in Table 2, our method’s average scores are higher than previous methods.
@@ 4.6 Ablation Study
Effect of Global Scan. We validate the effectiveness of the global scan, as listed in Table 3, the incorporation of global scan improves the overall performance of our method. For the multi head self attention module in global scan, the incorporation of multi-head self attention acquires improvement of Diversity and a degradation for FGD. Additionally, the ablation results demonstrate that incorporating Mamba enhances global scan’s capability to generate gestures with higher diversity and beat constancy. The FGD of generated gestures is better at the same time.
@@
Effect of Different Audio Encoder. To validate the effectiveness of the audio encoder, we replace the CNN-based audio encoder with pre-trained Whisper [47] and Vav2Vec2 [3], as listed in Table 4. Unlike CNN-based audio encoders that are randomly initialized and trained from scratch, when using Whisper and Wav2Vec2, we initialize the encoder using pre-trained weights and fix the parameters of the feature extractor. We observe a notable enhancement in facial generation when utilizing Whisper, however, the body generation results were subpar. In contrast, while Wav2Vec2 demonstrates some improvement in body generation, it results in a substantial decline in facial generation quality.
@@ A.2 Efficiency Analysis
Local Scan $0.00219\pm 0.0004$
Global Scan $0.00676\pm 0.0003$
@@ A.5 Pseudo Code
2:  $\mathbf{T}_{l-1}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{SelfAttention}(\mathbf{T}_{l-1})$
3:  $\mathbf{z}_{\mathbf{face}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{face}}(\mathbf{T}_{l-1}^{{}^{\prime}face})$
4:  $\mathbf{z}_{\mathbf{upperbody}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{upperbody}}(\mathbf{T}_{l-1}^{{}^{\prime}upperbody})$
5:  $\mathbf{z}_{\mathbf{lowerbody}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{lowerbody}}(\mathbf{T}_{l-1}^{{}^{\prime}lowerbody})$
6:  $\mathbf{z}_{\mathbf{hand}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{hand}}(\mathbf{T}_{l-1}^{{}^{\prime}hand})$
7:  /* process with different direction */
8:  for $o$ in {face, upperbody, lowerbody, hand} do
9:     $\mathbf{x}^{\prime}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SiLU}(\mathbf{Conv1d}_{o}(\mathbf{x}))$
10:     $\mathbf{B}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{B}}_{o}(\mathbf{x}^{\prime}_{o})$
11:     $\mathbf{C}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{C}}_{o}(\mathbf{x}^{\prime}_{o})$
12:     /* softplus ensures positive $\mathbf{\Delta}_{o}$/
13:     $\mathbf{\Delta}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}^{\mathbf{\Delta}}_{o}(\mathbf{x}^{\prime}_{o})+\mathbf{Parameter}^{\mathbf{\Delta}}_{o}))$
14:     /* shape of $\mathbf{Parameter}^{\mathbf{A}}_{o}$ is $(\mathtt{E},\mathtt{N})$/
15:     $\overline{\mathbf{A}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{Parameter}^{\mathbf{A}}_{o}$
16:     $\overline{\mathbf{B}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{B}_{o}$
17:     $\mathbf{y}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SSM}(\overline{\mathbf{A}}_{o},\overline{\mathbf{B}}_{o},\mathbf{C}_{o})(\mathbf{x}_{o}^{\prime})$
18:  end for
19:  /* get gated $\mathbf{y}_{o}$/
@@
20:  $\mathbf{y}_{\text{face}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{face}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{face}})$
@@
21:  $\mathbf{y}_{\text{upperbody}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{upperbody}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{upperbody}})$
@@
22:  $\mathbf{y}_{\text{lowerbody}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{lowerbody}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{lowerbody}})$
@@
23:  $\mathbf{y}_{\text{hand}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{hand}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{hand}})$
24:  /* residual connection */
@@
25:  $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}_{\mathbf{face}}^{\prime}+\mathbf{y}_{\mathbf{upperbody}}^{\prime}+\mathbf{y}_{\mathbf{lowerbody}}^{\prime}+\mathbf{y}_{\mathbf{hand}}^{\prime})+\mathbf{T}_{l-1}$
26:  Return: $\mathbf{T}_{l}$

### v4

# MambaTalk: Efficient Holistic Gesture Synthesis with Selective State Space Models

@@ 1 Introduction
State space models (SSMs) have recently shown significant potential in addressing challenges related to modeling sequences with low latency [12]. Inspired by continuous state space models from control systems and enhanced by HiPPO initialization [13], SSMs [16] show promise in addressing long-term forgetting issue. These advancements have been integrated into large-scale representation models [38, 41]. Some pioneering works have applied SSMs for tasks like language understanding [38, 41], content-based reasoning [12], and visual recognition [35, 57]. In our work, we further explore the potential of SSMs in co-speech gesture synthesis. We observe that directly applying the selective scan mechanism from Mamba [12] to gesture generation as a sequence modeling model would result in jittery outputs. To refine the generated gestures, we propose a two-stage modeling strategy. In the first training stage, we enhance the discrete motion priors derived from VQVAEs [53] by integrating velocity and acceleration losses. In the second stage, by utilizing motion priors from VQVAEs, we introduce individual learnable queries for different body parts, thereby alleviating the jittering issue. Meanwhile, considering that the direct application of Mamba encounters the challenge of limb movements across different body parts tending to average out, we propose a hybrid scanning approach in the second stage to enhance the motion representation in the latent space. Specifically, we refine the design of spatial and temporal modeling within latent spaces by introducing a global-to-local modeling strategy and integrating attention mechanisms along with a selective scanning approach into the framework’s design. Considering the significant differences in deformation and movement patterns among different body parts [32], we propose local and global scan modules for refining the latent space representations of the movements across various body parts. These approaches enable dynamic interaction and iterative refinement of different body parts while maintaining low latency, leading to more diverse and rhythmic gestures. Our contributions can be summarized as below:
@@ 3.1 Preliminaries
where $(\Delta A)^{-1}$ denotes the inverse of matrix $\Delta A$, $I$ denotes the identity matrix. The scan module within Mamba is designed to capture temporal patterns and dependencies across multiple time steps by applying a set of trainable parameters or operations to each segment of the input sequence. In our framework, Mamba serves as a sequence modeling tool for decoding gesture actions across different parts of the body. By modifying the decoder’s input and the range of features, we utilize Mamba to separately model the global motion features and local motion features of different body parts. These operations are learned during training and assist the model in processing sequential data.
@@ 3.3 Speech-Driven Selective State Spaces Gesture Synthesis Model
Speech Feature Extraction. For audio feature extraction, two CNN-based audio feature extraction networks are employed to respectively extract features from amplitude, raw audio and decoded tokens from Wav2vec2CTC [3]. Considering that the movements of body parts are not closely linked to raw audio, the audio encoder for body parts does not utilize raw audio as input. Specifically, we integrate these features along the channel dimension to obtain audio features $f_{A}=\{fa_{1},...,fa_{N}\}$. For processing speech input words, we employ pre-trained FastText [5] to obtain word embeddings, which are then refined by linear projections to produce text features $f_{T}=\{ft_{1},...,ft_{N}\}$. We further fuses features from the input modalities (e.g., audio and text features). The speaker ID embeddings $s_{id}$ are first combined with audio and text features through additive operation. By concatenating feature vectors along the channel dimension, we further apply linear transformations to determine the weight factors, and then integrating the features through an element-wise summation. The process can be formalized as:
@@
Global and Local Scans. Recognizing the diverse deformations and motion patterns in various body parts, we propose using global scan module and multiple local scan modules to model the movements of different body parts (e.g., face, hand, upper and lower body) with fused multi-modal features from previous modules. By acquiring the speech features from audio and text encoders, we first improve the perception of motion patterns among them using a global scan module by combining the speech features along the sequence dimension. Subsequently, by utilizing self-attention mechanism($\mathcal{F}_{\text{MHSA}}$), we model the global information across different sequence tokens. Following previous work [32], we establish a set of learnable parameters ($Q_{global}$) and employ masked motion to facilitate the acquisition of global information. Then, we employ self-attention mechanisms to enhance the global information and obtain the refined features. These refined features are fed into Mamba to extract temporal perceptual information. Considering the differences in representation between the body and face, we employ two independent MAMBA models to model the temporal features of the face and body, respectively. We then merge these features using a linear layer to obtain global features. The process can be fomulized as below:
@@
$\begin{gathered}\bar{{f}}_{global}=\mathcal{F}_{\text{MHSA}}(Q_{global}),\\ {f}_{speech}=\text{Mamba}([\bar{f}_{T},\bar{f}_{A}]),\\ \hat{f}_{global}=\text{Mamba}(\bar{{f}}_{global}),\\ {f}_{global}=\text{Linear}([{f}_{speech},\hat{{f}}_{global}]),\end{gathered}$ (7)
where [][][ ] denotes the concatenation operation of features in the dimension of the sequence. To enhance the generalization of the model, we incorporate the learnable queries to foster the queries’ ability to learn motion patterns. As shown in Figure 2, the queries from global scan are integrated with input speech features through a multihead cross-attention mechanism ($\mathcal{F}_{\text{MHCA}}$). This allows queries to learn the most relevant information from the speech input. The process can be formally defined as belows:
$f_{\text{refine}}=\mathcal{F}_{\text{MHCA}}(\bar{{f}}_{global},[\bar{f}_{T},\bar{f}_{A}]),$ (8)
where $f_{\text{refine}}$ denotes the feature of refined learnable queries. Utilizing the extracted perceptual features from various body parts, we proceed to employ Mamba to extract temporal features from the sequence, which can be formalized as belows:
${F}_{face}$ $=\text{Mamba}({f}_{refine}),$ (9)
where $F_{\text{face}}$ corresponds to the temporal features of facial motion. The same approach is utilized to generate temporal features for the hand, upper body, and lower body by inputting their respective perceptual features $f_{hand}$, $f_{upper}$, and $f_{lower}$ into the corresponding Mamba modules. One distinction is that we incorporate an additional self-attention layer before the Mamba layer to enhance the perception of body movements. The local latent features are then fed into their respective VQ-Decoders to produce the final motion predictions.
@@
Training Objectives. The model’s training objectives are composed of a composite loss function that harmonizes reconstruction and cross-entropy losses. This design aims to augment the accuracy of motion generation, encompassing the face, hands, upper, and lower body. The loss of latent reconstruction, represented by $L_{reclatent}$, is quantified using the Mean Squared Loss (MSELoss). Here, $z_{i}$ corresponds to the true latent vectors, while $\hat{z}_{i}$ are the vectors reconstructed by the model. The latent reconstruction loss is expressed as:
@@
where $N$ denotes the number of frames. Concurrently, to encourage diversity in the generated motions, we optimize the cross-entropy loss for latent code class classification $L_{cls}$. Specifically, we employ Negative Log-Likelihood Loss (NLLLoss), where $y_{i}$ represents the true class labels for each sample, and $\hat{y}_{i}$ denotes the model’s predicted class labels. This loss is calculated as the negative sum of the logarithm of the predicted probabilities for the correct classes:
@@ 4.1 Experiments Setting
We train and evaluate on the BEAT2 dataset proposed by [32]. BEAT2 contains 60 hours of data with high finger quality for 25 speakers (12 female and 13 male). The dataset comprises 1762 sequences, each with an average duration of 65.66 seconds. Each sequence includes a response to a daily inquiry. We split datasets into 85%/7.5%/7.5% for the train/val/test set. We follow previous work [32] to select data from Speaker 2 for training and validation to ensure fair comparison.
@@ 4.4 Quantitative Results
As shown in Table 1, our method attains the lowest FGD and highest BC when compared to the previously top-performing method. This highlights the superior ability of MambaTalk in discerning and correlating the audio-motion beats. The lowest FGD also emphasizes the high quality and naturalness of our generated movements, showing the ability of MambaTalk to capture real motion dynamics. This also demonstrates the authenticity of our generated motions, affirming the successful capture of inherent motion characteristics. Some results are marked as “-” because these methods can not generate facial movements. Moreover, our method outperforms previous methods in terms of MSE and LVD, with substantial improvements of 18.11% and 8.72%, respectively. These two enhancements highlight the superior accuracy and fidelity of our method in capturing fine-grained details, affirming its efficacy in synthesizing realistic and authentic facial motions.
Table 1: Quantitative results on BEAT2. FGD (Frechet Gesture Distance) multiplied by $10^{-1}$, BC (Beat Constancy) multiplied by $10^{-1}$, Diversity, MSE (Mean Squared Error) multiplied by $10^{-7}$, and LVD (Learned Vector Distance) multiplied by $10^{-5}$. The best results are in bold.
@@
MambaTalk (Ours) - 5.366 $\mathbf{7.812}$ 13.05 $\mathbf{6.289}$ $\mathbf{6.897}$
@@ 4.5 Qualitative Analysis
User Study. We conducte a user study to assess the visual quality of the generated co-speech 3D gestures. For each method under comparison, we produce 10 gesture samples, which were then converted into video clips for evaluation by 39 participants. In each evaluation session, participants are presented with 20 seconds video clips generated by various models. They are instructed to assess the clips across the following dimensions: (i) naturalness, (ii) appropriateness, (iii) synchrony and (iv) smoothness. For naturalness, they evaluate the similarity of the generated gestures to those made by humans, paying attention to the authenticity and smoothness of the movements. In terms of appropriateness, they consider the alignment of the gestures with the spoken content, taking into account both the explicit meaning and the underlying semantics. For synchrony assessment, they examine the timing of the gestures in relation to the speech rhythm, audio, and facial expressions to ensure a harmonious and integrated performance. For smoothness, they assess the gestures for any abrupt stops or unnatural jerks that might indicate a lack of fluidity in motion. We mainly compare two state-of-art methods with our proposed method (with and without VQVAE): CaMN [33], EMAGE [32], and the ground truth. As presented in Table 2, our method’s average scores are higher than previous methods.
@@ 4.6 Ablation Study
Effect of Global Scan. We validate the effectiveness of the global scan, as listed in Table 3, the incorporation of global scan improves the overall performance of our method. For the multi-head self-attention module in global scan, the incorporation of multi-head self-attention acquires improvement of Diversity and a degradation for FGD. Additionally, the ablation results demonstrate that incorporating Mamba enhances the global scan’s capability to generate gestures with higher diversity and beat constancy. The FGD of generated gestures is better at the same time.
@@
Effect of Different Audio Encoder. To validate the effectiveness of the audio encoder, we replace the CNN-based audio encoder with a pre-trained Whisper [47] and Vav2Vec2 [3], as listed in Table 4. Unlike CNN-based audio encoders that are randomly initialized and trained from scratch, when using Whisper and Wav2Vec2, we initialize the encoder using pre-trained weights and fix the parameters of the feature extractor. We observe a notable enhancement in facial generation when utilizing Whisper, however, the body generation results were subpar. In contrast, while Wav2Vec2 demonstrates some improvement in body generation, it results in a substantial decline in facial generation quality.
@@ A.2 Efficiency Analysis
Global Scan $0.00219\pm 0.0004$
Local Scan $0.00676\pm 0.0003$
@@ A.5 Pseudo Code
2:  $\mathbf{z}_{\mathbf{face}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{face}}(\mathbf{T}_{l-1}^{{}^{\prime}face})$
3:  $\mathbf{z}_{\mathbf{upperbody}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{upperbody}}(\mathbf{SelfAttn}(\mathbf{T}_{l-1}^{{}^{\prime}upperbody}))$
4:  $\mathbf{z}_{\mathbf{lowerbody}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{lowerbody}}(\mathbf{SelfAttn}(\mathbf{T}_{l-1}^{{}^{\prime}lowerbody}))$
5:  $\mathbf{z}_{\mathbf{hand}}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{hand}}(\mathbf{SelfAttn}(\mathbf{T}_{l-1}^{{}^{\prime}hand}))$
6:  /* process with different parts of human body */
7:  for $o$ in {face, upperbody, lowerbody, hand} do
8:     $\mathbf{x}^{\prime}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SiLU}(\mathbf{Conv1d}_{o}(\mathbf{x}))$
9:     $\mathbf{B}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{B}}_{o}(\mathbf{x}^{\prime}_{o})$
10:     $\mathbf{C}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{C}}_{o}(\mathbf{x}^{\prime}_{o})$
11:     /* softplus ensures positive $\mathbf{\Delta}_{o}$/
12:     $\mathbf{\Delta}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}^{\mathbf{\Delta}}_{o}(\mathbf{x}^{\prime}_{o})+\mathbf{Parameter}^{\mathbf{\Delta}}_{o}))$
13:     /* shape of $\mathbf{Parameter}^{\mathbf{A}}_{o}$ is $(\mathtt{E},\mathtt{N})$/
14:     $\overline{\mathbf{A}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{Parameter}^{\mathbf{A}}_{o}$
15:     $\overline{\mathbf{B}}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}_{o}\bigotimes\mathbf{B}_{o}$
16:     $\mathbf{y}_{o}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{SSM}(\overline{\mathbf{A}}_{o},\overline{\mathbf{B}}_{o},\mathbf{C}_{o})(\mathbf{x}_{o}^{\prime})$
17:  end for
18:  /* get gated $\mathbf{y}_{o}$/
@@
19:  $\mathbf{y}_{\text{face}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{face}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{face}})$
@@
20:  $\mathbf{y}_{\text{upperbody}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{upperbody}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{upperbody}})$
@@
21:  $\mathbf{y}_{\text{lowerbody}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{lowerbody}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{lowerbody}})$
@@
22:  $\mathbf{y}_{\text{hand}}^{\prime}$ : $(\mathtt{B},\mathtt{M},\mathtt{E})$leftarrow $\mathbf{y}_{\text{hand}}\bigodot\mathbf{SiLU}(\mathbf{z}_{\mathbf{hand}})$
23:  /* residual connection */
@@
24:  $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{M},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}_{\mathbf{face}}^{\prime}+\mathbf{y}_{\mathbf{upperbody}}^{\prime}+\mathbf{y}_{\mathbf{lowerbody}}^{\prime}+\mathbf{y}_{\mathbf{hand}}^{\prime})+\mathbf{T}_{l-1}$
25:  Return: $\mathbf{T}_{l}$